Joanna is 'x' years old. Compared to Joanna, Betty is 6 years older and Mary is 2 years older. What is the total of their ages?
Answer by macston(5194)

.
J=Joanna's age; B=Betty's age=J+6yrs; M=Mary's age=J+2yrs
J+B+M=J+(J+6yrs)+(J+2yrs)=3J+8yrs
J=x
Sum of ages=3J+8yrs Substitute for J.
Sum of ages=3x+8 yrs